OpenStudy (anonymous):

. A 110-foot long piece of rope is cut into two pieces so that one piece is two ninths as long as the other piece. Find the length of each piece

OpenStudy (anonymous):

20,90

OpenStudy (anonymous):

So let the length of piece one be x. Then the length of piece two is (2/9)x We know the two pieces added must equal the original 110 feet. x+(2/9)x=110 (11/9)x=110 x=90

OpenStudy (anonymous):

x+(2/9)x=110 11x/9=110 x=90 theother piece=(2/9)*90=20